please solve
18y4-42y3

4y2+25y+25

54t3+16v3

y2-3y-10=0

Answer by stanbon(48579)

please solve
18y^4-42y^3
= 6y^3(3y-7)
----------------------
4y^2+25y+25
=4y^2+20y+5y+25
=4y(y+5)+5(y+5)
=(y+5)(4y+5)
----------------
54t3+16v3
2(27t^3+8v^3)
=2(3t+2v)((3t)^2-(3t)(2v)+(2v^2))
=2(3t+2v)(9t^2-6tv+4v^2)
---------------
y^2-3y-10=0
(y-5)(y+2)=0
y=5 or y=-2
--------------
Cheers,
Stan H.